OpenStudy (gabylovesyou):

Shades R Us charges $20 per day to rent a lounge chair and $15 per day to rent an umbrella. Dan and Lisa paid a total of $245 to rent a lounge chair and an umbrella each during their vacation. Lisa rented the chair and umbrella for 1 day less than Dan. The following equation represents this situation, where x represents the number of days Dan rented the lounge chair and umbrella. 20x + 15x + 20(x − 1) + 15(x − 1) = 245 What is the total amount Dan paid to rent the lounge chair and umbrella during his vacation? $80 $105 $140 $175

OpenStudy (gabylovesyou):

is it 105 ?

OpenStudy (ashleyisakitty):

If he spends 35 dollars a day for four days he would get...

OpenStudy (gabylovesyou):

what am i supposed to do ? solve for x ?

OpenStudy (ashleyisakitty):

This is what i did... so basically you break it down like so: 20x + 15x + 20(x - 1) + 15(x - 1) = 245 35x + 20(x - 1) + 15(x - 1) = 245 35x + 35(x-1) = 245 35x + 35x - 35 = 245 70x - 35 = 245 70x = 280 x = 4 then it says that he pays 35 dollars a day for the lounge and umbrella. therefore it would be 4(35). What is the answer?

OpenStudy (gabylovesyou):

140 :) thank you soo much

OpenStudy (ashleyisakitty):

:~) no problem!